How much pure gold should be melted into 5 grams of a 62% gold alloy to make it an 85% gold alloy?
Answer by stanbon(75887)

How much pure gold should be melted into 5 grams of a 62% gold alloy to make it an 85% gold alloy?
------
Equation:
gold + gold = gold
0.62*5 + 1.00x = 0.85*(5+x)
Multiply thru by 100:
62*5 + 100x = 85*5+85x
15x = 23*5
x = (1/3)23
x = 7 2/3 oz (amt. of pure gold needed)
===========================================
Cheers,
Stan H.